40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 171 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  8995   Mon Aug 12 12:57:59 2013 JenneUpdateCDSX-End Green ASS - Roundup

Quote:
  1. The SIMULINK model has been modified to accommodate an option to dither the cavity mirrors and not the PZT mirrors. Details are as follows:
    • I have sent the LO signals (CLK,SIN and COS) from the ASS model to the ASX model via the RFM model. Appropriate changes were made to all these three models, and recompiling and restarting the models was done without issue. The SIN and COS signals are used to demodulate green transmission at the dither frequencies. ***The CLK signal is not required to be sent between models as it is not being used by ASX (I turn the dither ON using the channels already set up for ASS). I realised this a little late, and at present the ASS and RFM models are compiled such that the CLK signal is also sent from ASS to RFM. This can be removed, thus freeing up 4 unnecessary inter-process communication channels. Also, I am not too sure if this is relevant, but the maximum computation time of both the RFM and ASX models seem to have gone up after I added these inter-process communication links.***

    •  

 Getting rid of the LO transmission will certainly help / be good.  After adding these channels, the RFM model is regularly hitting 62usec (out of a max acceptable of 60).

I'm not really sure why the ASS was involved in this.  I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner.  Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model.  I had thought that this was expressly forbidden (although I'm not sure why).  I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level.

  8994   Mon Aug 12 10:44:22 2013 ranaUpdateLSCPRMI(sb) lock recovered

  In the past, we used to use Stefan's 'ezcademod' or Matt's 'ezlockin' to do auto phase adjustment.

JoeB / Jamie are working on python replacements for these tools, but in the near term possibly I can make a bash script to use ezcaservo and the existing LOCKINs to do this.

  8993   Sat Aug 10 05:53:51 2013 gautamUpdateCDSX-End Green ASS - Roundup

Over the last three days, I've had the interferometer to test and optimize the ASX Servo. Based on what I have seen, I think the conclusion is that with the current parameters, the servo does its job provided the input pointing set up at the endtable with the coarse adjustment knobs is reasonably good. Once the cavity is aligned and IR transmission maximized using ASS, I have been able to get the green transmission up to 0.8 which is close to the best we had pre-vent. I have not been elogging regularly over the last few days, so this one is going to be a longish one.


Major changes made:

  1. The SIMULINK model has been modified to accommodate an option to dither the cavity mirrors and not the PZT mirrors. Details are as follows:
    • I have sent the LO signals (CLK,SIN and COS) from the ASS model to the ASX model via the RFM model. Appropriate changes were made to all these three models, and recompiling and restarting the models was done without issue. The SIN and COS signals are used to demodulate green transmission at the dither frequencies. ***The CLK signal is not required to be sent between models as it is not being used by ASX (I turn the dither ON using the channels already set up for ASS). I realised this a little late, and at present the ASS and RFM models are compiled such that the CLK signal is also sent from ASS to RFM. This can be removed, thus freeing up 4 unnecessary inter-process communication channels. Also, I am not too sure if this is relevant, but the maximum computation time of both the RFM and ASX models seem to have gone up after I added these inter-process communication links.***
    • The rest of this part of the servo is a replica of the part where PZT mirrors are dithered. At present the servo output is the sum of its two branches (PZT mirror dither branch and cavity mirror dither branch) which works fine under the assumption that at any one time, only one arm will run. Ideally, the summing block should be replaced by a switch. However, when I tried (in an earlier attempt to include the cavity dither) to do this and restart the model, c1iscex crashed, and so I decided against using the switch block for this trial.
    • The control signal generated using green transmission demodulated at the ETM dither frequencies are used to actuate on M1 while the ITM ones are used to actuate on M2. Of course, by setting the appropriate off-diagonal elements in the output matrix, this can be modified as desired.
  2. The main MEDM screen has been updated to reflect the new additions to the SIMULINK model. Screenshot is attached. The picture isn't entirely accurate as the monitor channels in the upper row actually show the servo output + slider output. This needs to be changed in the model, and a new set of monitors need to be added to the MEDM screen. In the end, we require four sets of monitor-points in the model: PZT dither servo output, cavity dither servo output, sum of these with any offset from the PZT sliders, and the sum of the latter with the dither signal (this is what eventually goes to the PZT mirrors while the dither is on).
  3. I added scripts to the MEDM screen that turn the PZT mirror dither servo on and off. Note that when you want to run a new script on an MEDM screen using medmrun, you need to change the permissions of the file by going to the path where your script is located and running chmod 755 <name of script>. Manasa has updated the same on the wiki.

 Details of tests runs:

For the most part, I have been trying to optimize the PZT mirror dither servo. To this end, I did the following:

  • Went to the X-end and fixed the input pointing which was not optimal. Manasa first aligned the arm and ran ASS to maximize the IR transmission. I then used the coarse adjustment knobs on the mirror mounts to get the green transmission up to ~0.6.
  • I then set the following parameters in the servo (these are all in the script, path to which is /opt/rtcds/caltech/c1/scripts/ASX):
    1. LO frequencies of 10, 19, 34 and 39 Hz respectively for M1 PIT, M1 YAW, M2 PIT and M2 YAW.
    2. LO amplitudes of 75 for all the four degrees of freedom (determined by using PZT calibration to see what amplitude would couple 10% of power into the first higher-order-mode assuming a perfectly aligned beam to start with.
    3. LIA BP filters centered at the above frequencies with 2Hz passband on either side.
    4. LIA LP filters with corner frequency 0.5 Hz.
    5. LIA Signal filter bank gain set to 100 for all degrees of freedom.
    6. LIA Demod I phase filter bank gain set to 5 for all degrees of freedom.
    7. Control filter gains to 1 for all degrees of freedom (control filters are all integrators).
    8. Demod phase set to 0 for all degrees of freedom. I did not really try to optimize this but the servo seems to be doing reasonably well even with this setting.
    9. Overall servo gain to 1 (the servo worked well when I increased this to 5 as well, but became unstable when I increased it further).
  • I ran the servo. Observations were as follows:
    • Having fixed the input pointing to get green transmission up to ~0.6, the servo was able to improve it to ~0.8, which is the best we had after hours spent at the X-end prior to the vent.
    • Given a good input pointing, we can use the PZT mirrors to lock to 00 mode from some misaligned state using either the sliders, or by leaving the servo on, and helping it out at the points where it gets 'stuck' in some higher mode using either the sliders or by toggling the shutter.
    • In order to recover green transmission of ~0.8, it was often necessary to first run ASS to optimize the IR transmission. Otherwise, green-transmission saturates at ~0.6 or 0.4 depending on the misalignment of the arm cavity mirrors. The servo was unable to change the input pointing enough to deal with overly misaligned cavity mirrors. 
    • The servo is sometimes capable of bringing about mode-hopping from a higher order mode to a lower one, though this is not always the case as the PDH lock is sometimes too strong, in which case I toggled the shutter after which the servo kicked in.
    • I tested the servo under as many different conditions as I could. For instance, having left the green shutter open overnight, I saw that the transmission had fallen from 0.8 (which was what we saw on Thursday night) to ~0.4 on Friday morning. Running the servo got the transmission up to 0.6. I then asked Manasa to run ASS, (while leaving the ASX servo on), after which point the green transmission went up to 0.8. Sometimes, the servo locks to a 'bad' 00 mdoe, where the transmission saturates at ~0.2, but toggling the shutter fixes this most of the time.

Attempt to measure transfer function:

One of the things that came up during my presentation was how fast the loop was capable of responding. I was able to get a quantitative idea of this by playing around with the overall servo gain. Initially, it took ~30 seconds for the servo to get the transmission up to its peak value, with a servo gain of 1. When I ramped this up to 5, the response was much faster, with the peak transmission being reached in ~5seconds. 

 

I wanted to get a more quantitative picture, and hence tried to measure the transfer function by first injecting an excitation into the 'SIG' filter-bank in the demodulation stage. However, coherence between the IN1 and IN2 signals was very poor for all the amplitude configurations I tried. At Jenne's suggestion, I tried injecting the excitation at the control-filters stage, but found no improvement. Perhaps the amplitude envelope was wrong or the measurement technique has to be rethought. 


 Misc remarks:

  1. M1 is the first steering mirror and M2 is the second one (right before the beam enters the arm cavity).
  2. Though I have added the cavity dither feature to the model, I was not able to optimize this servo. Some calculations need to be done to get an estimate of the output matrix, after which the filter gains etc can be optimized.
  3. Today, I cleaned up my temporary setup at the SP table to calibrate the PZTs. Most of the hardware for the Y-end is now in the tupperware box. The QPD and laser have been restored to the optical bench next to MC2 where I found them. The second KEPCO HV supply which I had set up has now been installed at 1Y4 in anticipation of the PZT mirrors at the Y-endtables. It is currently powered OFF.
  4. Performance plots to follow as I have not pulled the data out yet.
  5. I had bought a cake from chandler today in an effort to clear my meal plan, but in the rush in the afternoon, completely forgot about it. It is in the fridge, and is strawberry tart, hope it tastes good.

 


 New MEDM screen:

New_ASX_MEDM_MAIN.pdf 

  8992   Fri Aug 9 22:51:37 2013 KojiUpdateLSCPRMI(sb) lock recovered

PRMI(sb) lock was recovered


PRMI lock

- Stared at the time series data of the REFL demod signals, and decided to use REFL165I&Q for the locking.

- Jiggled the demodulation phase of REFL165 and POP110. Changed the servo gains.

- Finally found a short lock. Further optimized the parameters.

- PRM ASC was turned on by giving the identity matrices for the input and output matrices.
  Now just hitting the up button is sufficient to engage the ASC servo.

- Under the presence of the ASC, the PRMI is indefinitely locked as before.

- Reacquisition is also instantaneous. (It acquires even if the ASC is left "on".)

- Actually the lock is somewhat robust even when the PRM ASC is not used.
  This is VERY GOOD as we can skip one of the steps necessary for the full lock.

  Although, the seismic on Friday night is very quiet.
  The spot motion at POP seems to be somewhat pitch/yaw mixed, in stead of previous "totally-dominated-by-yaw" situation.

- We are ready to implement ASS for PRM

Demod phase adjustment

- Shook PRM at 580Hz / 100cnt

- Swept the demod phase of REFL165 such that the PRM peak is minimized in the Q signal

- Open DTT. Measured transfer functions between REFL165I and the Q signals of each PD.

- Minimized the PRCL signal coupling in the signals.

- The resolution of the adjustment was ~1deg.

Locking test with PRM/BS

Tried the lock acquisition only with PRM and BS. (cf. http://nodus.ligo.caltech.edu:8080/40m/8816)

This just worked nicely.


Today's locking parameters:

PRMI(sb) lock:

MC Trans: 17500
POP110I (in lock): 150

PRCL Source: REFL165(I) 106deg / 45dB / Normalization SQRT(10 POP110I) / Input MTRX 1.0
PRCL Trigger: POP110I x 1.0 50up 25down
PRCL Servo: G=+3.5 Acq: FM4/FM5 Opr: FM2/FM3/FM6/FM7
PRCL Actuator: PRM +1.0

MICH Source: REFL165(Q) 106deg / 45dB / Normalization SQRT(0.1 POP110I) / Input MTRX 1.0
MICH Trigger: POP110I x 1.0 50up 25down
MICH Servo: G=-10 Acq: FM4/FM5 Opr: FM2/FM3/FM6
MICH Actuator: (ITMX -1.0 / ITMY +1.0) or (BS 0.5 / PRM -0.267)

Demod phases:

AS55 -17deg
REFL11 135deg
REFL33 -18deg
REFL55 120deg
REFL165 106deg

  8991   Fri Aug 9 21:05:28 2013 KojiUpdateSUSfixed: SRM coils fine - problem with slow bias slider

Now the SRM Yaw bias in yaw is functional without any strage behavior.
The problem was found at the connector of the flat ribbon cable from the DAC to the cross connect.

I used the extender board to diagnose the SRM coil driver circuit at 1X4.
The UL coil input did not show any sign of voltage no matter how the bias slider was jiggled.

I opened the side panel of the rack and found the signal was absent at the cross connect which relays two flat ribbon cables
for the SRM coil driver. I checked the DAC output with a multimeter. All the bias outputs were OK at the DAC.

Then I opened the IDC connector at the DAC side of the crossconnect as the signal was already missing there.
I found that the flat ribbon cable was a half line shifted from the supposed location.
This resulted a short circuit of the DAQ +/- pins for the SRM UL coil.

I recrimped the connector and now the SRM Yaw slider is back.
This changed the nominal position of the SRM. The new slider values were saved.

  8990   Fri Aug 9 16:49:35 2013 Jenne, manasaUpdateElectronicsPost-vent alignment cont'd - RFPDs

Notes to the fiber team:

I am aligning beam onto the RFPDs (I have finished all 4 REFL diodes, and AS55), in preparation for locking. 

In doing so, I have noticed that the fiber lasers for the RFPD testing are always illuminating the photodiodes!  This seems bad!  Ack!  

For now, I blocked the laser light coming from the fiber, did my alignment, then removed my blocks.  The exception is REFL55, which I have left an aluminum beam dump, so that we can use REFL55 for PRM-ITMY locking, so I can align the POP diodes.

EDIT:  I have also aligned POP QPD, and POP110/22.  The fiber launcher for POP110 was not tight in its mount, so when I went to put a beam block in front of it and touched the mount, the whole thing spun a little bit.  Now the fiber to POP110 is totally misaligned, and should be realigned.

What was done for the alignment:

1. Aligned the arms (ran ASS).

2. Aligned the beam to all the REFL and AS PDs. 

3. Misaligned the ETMs and ITMX. 

4. Locked PRM+ITMY using REFL11.
The following were modified to enable locking
(1) PRCL gain changed from +2.0 to -12.
(2) Power normalization matrix for PRCL changed from +10.0 to 0.
(3) FM3 in PRCL servo filter module was turned OFF.

5. POP PDs were aligned.

  8989   Thu Aug 8 21:25:36 2013 KojiUpdateGeneralPost-vent alignment cont'd

- IPANG aligned on the QPD. The beam seems to be partially clipped in the chamber.

- Oplev of the IFO mirrors are aligned.

- After the oplev alignment, ITMX Yaw oplev servo started to oscillate. Reduced the gain from -50 to -20.

  8988   Thu Aug 8 18:47:41 2013 SujanSummaryPEMUsing weiner filters for subtracting signals MC_L and GUR2_X

I used MC_L signal from the Mode Cleaner as the desired signal with GUR2_X as witness signals. I observed good subtraction where coherence is high. But there was noise added in other frequency bands. I am not sure how to avoid that.

Please find attached documents that contains relevant plots.

  8987   Thu Aug 8 18:43:12 2013 SujanSummaryPEMOptimally subtracting signals from two seismometers.

An exercise of optimally subtracting one seismometer signal by another using weiner filters was done. Results have been summarized document attached.

  8986   Thu Aug 8 11:18:46 2013 manasaUpdateComputer Scripts / ProgramsUnused scripts in ASS moved

I was receiving missing path error when I was trying to measure the MC spot positions. Jenne pointed out that Koji had moved all the unused scripts in scripts/ASS to /scripts/ASS/OBSOLETE yesterday and in the process one of the scripts that the MC spot position measurement script calls for (MeasureSpotPositions.py) must have also been moved to the OBSOLETE directory. I moved the script to /scripts/ASS/MC so that we know the script is being used and also changed its path in the main script.

  8985   Thu Aug 8 10:31:28 2013 SteveUpdateVACfew reminders of this vent

 1, Vacuum envelope grounds must be connected all times!  After door removal reconnect both cables immediately.

 2, The crane folding had a new issue of getting cut as picture shows.

 3, Too much oplev light is scattered. This picture was taken just before we put on the heavy door.

 4, We were unprepared to hold the smaller side chamber door 29" od of the IOC

 5, Silicon bronze 1/2-13 nuts for chamber doors will be replaced. They are not smooth turning.

 

  8984   Thu Aug 8 07:57:23 2013 SteveUpdateVACpumpdown at 48 hrs

Quote:

 

 TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.

 

  8983   Wed Aug 7 23:40:49 2013 gautamUpdateCDSX-End Green ASS - A first update

 I have done some preliminary testing of the X-End Green ASS Servo. I will write a more detailed elog about this soon, but I thought I'd note down the important stuff here.


Yesterday, while we were venting, I aligned the X-arm to the green using the sliders on IFOalign, maximizing the transmission. Then I retook a power spectrum so as to determine the LO frequencies. Jenne pointed out that LO frequencies should not be integers (it usually suffices to append a .098725 or something to the frequency) so I made the necessary changes.

I did a first run of the servo yesterday, and more runs today. Notable points:

  1. I was able to lock to 00 from a 08 or 09 mode using the PZT sliders
  2. The green transmission having locked to 00 was ~0.2. I then ran the servo and got it up to ~0.4 and then 0.6 (see time series plot attached). The servo was able to recover this level of transmission after misaligning the steering mirrors using the PZT sliders.
  3. This was not the optimal transmission level as when Koji moved ETMX a little, the transmission improved.
  4. The actuators are degenerate. Most of the time, only two of the four servos are doing anything significant. This is probably because of the fact that the two steering mirrors are so close to each other, that moving one or the other produces virtually the same effect. I do however have some cool videos of mode-hopping :)
  5. The range of actuation of the PZTs is probably not enough to maximize the green transmission from an arbitrary state because of point 4 (i.e. we need to move one mirror in some direction a lot, and move the other a lot to compensate for the change, and the overall gain in input pointing/alignment is marginal). It may be that things will be slightly better at the Y-end. It would also be interesting to see if there is any improvement in the servo performance by dithering the cavity mirrors as opposed to the PZT mirrors.
  6. To this end, I tried modifying the c1asx model to incorporate an option to dither the cavity mirrors. The plan was to make a second set of LOs in the model that output to ITMX and ETMX suspensions. However, for some reason, when I recompiled the model and restarted it, c1iscex crashed. Parity has now been restored. Note that in order to accommodate the new LOs, I had to make changes to C1SUS, C1RFM and C1SCX as well. I have since removed all my additions, saved, built and installed these models, but have not restarted them (with the exception of C1SCX which restarted when I manually restarted c1iscex). 
  7. The plan tomorrow is to try incorporating cavity dither into the model again. This time, I'll try grabbing the LO-related signals from c1ass directly, as I am not clear why my approach did not work.

More details to follow.

time-series.pdf

  8982   Wed Aug 7 22:18:43 2013 KojiUpdateASCASS update

While Gautam is working on the Xarm green ASS...

The EPICS monitor points for the ASS actuators were added to the ASS model.

This will be used for the offloading the ASS actuations to the alignment biases.
As this modification allowed us to monitor the actuation apart from the dithering,
now we can migrate the ASS actuation to the fast alignment offset on the suspension.
This modification to the offset moving scripts were also done.

Screenshot-Untitled_Window.png

  8981   Wed Aug 7 21:52:11 2013 JenneUpdateSUSSRM coils fine - problem with slow bias slider

[Koji, Jenne]

We have looked a little more at the SRM situation.  We aligned the SRM, and then aligned the oplev, so that we had a convenient monitor of the optic's motion.

When we use the _COMM channels, which are the usual ones on the IFO_ALIGN screen, the pitch slider makes pitch motion, but the yaw slider makes the oplev spot move ~45degrees from horizontal.

However, when we use the bias channels that are in the front end model, parallel to the ASC path, pitch moves pitch, and yaw moves pure yaw.

So, we conclude that the SRM coils are fine, and there is something funny going on with the slow part of the actuation. 

Koji restarted the slow computer susaux, and burt restored it, but that did not fix the situation.  We went inside and looked at all of the ribbon cable connections, and pushed them all in, but that also has not fixed things. 

We have been looking at D010001-b, the coil driver board, and we think that's where the summing resistor network between the slow bias slider, and the coil outputs from the fast model exists.  (It's not 100% clear, but we're confident that that's what is going on). 

Tomorrow, we will pull the SRM's coil driver board, and see if any of the components in the slow slider path, before the summing point, look burned / broken / bad.

  8980   Wed Aug 7 19:16:20 2013 JenneUpdateASCASS setting up accelerated more

I have furthered Koji's work, and moved the filter on/off state for all the filter banks also to the burt snapshot.  

Turning on the ASS is now much faster than it was originally, with the ezcawrites in series.

  8979   Wed Aug 7 15:51:53 2013 Alex ColeConfigurationElectronicsRF Switch Change

For the photodetector frequency response project, our new RF Switch Chassis (NI pxie-1071) arrived today. I took the switches out of the old chassis (Note for future generations: you have to yank pretty darn hard) and put them in the new chassis, which I mounted in rack 1Y1 as pictured. 

The point of this new chassis is that its controller is compatible with our control room computer setup. We will be able to switch the chassis using TCP/IP or telnet, aiding in our automation of the measurement of photodetector frequency response.

  8978   Wed Aug 7 15:36:29 2013 SteveUpdateVACTP2 drypump replaced

 

 TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.

Where are the foreline pressure gauges? These values are not on the vac.medm screen.

The new tip seal dry pump lowered the small turbo foreline pressure 10x

TP2fl after 2 day of pumping 65mTorr

  8977   Wed Aug 7 15:32:37 2013 KojiUpdateASCASS setting up accelerated (slightly)

I moved bunch of ezcawrite from the ASS Dither On script to a snapshot file.

This accelerated a half of the "up" time but still switching part is not in the snapshot.

If you find anything wrong with ASS, please notify me.

  8976   Wed Aug 7 10:51:49 2013 SteveUpdateVACPumpdown at 26hrs

 Cold cathode gauge CC1 -h (horizontal) just coming on 9.2e-5 Torr

P2 is the fore line pressure of the maglev. One can see the 4 Torr load during switching over to turbo pumping.

CC4  5e-9 Torr is the performance of the maglev pumping on the RGA only.

The annuloses are not pumped  now.  They are closed off to see how much outgassing plus leak they have.

Configuration: vacuum normal, annuloses not  pumped

Condition: normal

Precondition: 14 days at atm, IOO chamber north door was taken off as a new entrance, the ETMX chamber was not opened.

What is new in the vacuum system: new P1 pirani gauge, gold plated clean allen wrench and ..........what else was dropped?

Note: the wireless laptop did not fail once all day yesterday. I want to give credit to the person who is responsible for this.

 

  8975   Wed Aug 7 10:09:30 2013 KojiUpdateLSCArms locked in IR, aligned. IFO at nominal power

I have a concern about the SRM suspension. The yaw alignment bias produces huge pitch coupling.

This could be a connector issue or the rubbing of the mirror on the EQ stops.

We have the photos of the magnets and they were not touching the OSEMs.

  8974   Tue Aug 6 19:53:15 2013 JenneUpdateLSCArms locked in IR, aligned. IFO at nominal power

[Koji, Manasa, Jenne]

The Y arm was locked in IR, and we saw flashing in the Xarm (Gautam had the Xarm for green work when we began).  I checked IPANG, and the beam was beautifully unclipped, almost perfectly centered on the first out of vacuum mirror.  I aligned the beam onto the QPD.

We then swapped out the MC Y1 that we use at low power, and replace the usual 10% BS, so that we wouldn't crispy-fry MC REFL.  Manasa adjusted the half wave plate after the laser, to maximize the power going toward the PMC.  We relocked the PMC, and see transmission of ~0.84, which is at the high side of what we usually get.  The beam was aligned onto MC REFL and centered on the WFS, and the MC was locked at nominal power.  Koji tweaked up the alignment of the MC, and ran the WFS offset script.  I aligned beam onto POP QPD and POP110 coarsely (using a flashing PRC, not a locked PRM-ITMY cavity, so the alignment should be rechecked).  The arms have both been locked and aligned in IR....the green beams need to be steered to match the current cavity axis. 

The AS beam, as well as REFL and POP, are all coming out of the vacuum nicely unclipped. 

Notes:  When Koji was aligning the SRM to get the SRC cavity roughly aligned (the AS flashes all overlapping), we noticed that there is some major pitch-yaw coupling.  Serious enough that we should be concerned that perhaps some connector is loose, or an actuator isn't working properly.  This should be checked.

Moral of the story:  Coarse alignment of all mirrors is complete after pump-down and we have IR locked and aligned to both arms at nominal power.

 

Still to do:

* Restore PRM, align beam onto the REFL PDs. 

* Lock PRM-ITMY cavity, align beam onto POP PDs.

* Align AS beam onto AS55. 

* Recenter all oplevs.

* Recenter IPPOS and IPPANG at nominal power.

* Start locking!!

  8973   Tue Aug 6 16:45:31 2013 SteveUpdateVACPumpdown is completed

 IFO  P1=1mT PSL shutter is opened

 

  8972   Tue Aug 6 16:36:51 2013 gautamUpdateCDSChoosing LO Amplitudes and Frequencies-revised

I redid the power spectrum measurement for the X-arm green transmission after aligning the arm to green using the ITMX/ETMX Pitch and Yaw sliders on IFOalign.

The Y-axis now reflects the relative intensity noise (RIN), which I obtained by taking the average value of the X-arm green transmission using tdsavg. Based on this measurement, I have now picked four new frequencies at which to try and modulate the PZT mirrors: 10, 19, 34 and 39 Hz. Bandpass filters in the LIA stage have been appropriately modified. 

Power Spectrum:

powerSpec0806.pdf

  8971   Tue Aug 6 12:43:23 2013 Alex ColeConfigurationElectronicsAS Table and Rack 1Y1 Additions

For the photodetector frequency response project, I finished the construction of our baluns chassis and mounted it in rack 1Y1 (1st picture).

After consulting with Jenne, I mounted the fiber launcher for REFL165 on the AS table such that it would not cause an obstruction. I aligned the launcher using a multimeter to monitor the DC output of REFL165, but looking at the data I got, it seems I need to do a better alignment/focusing job to get rid of a bunch of noise.

  8970   Tue Aug 6 10:06:29 2013 SteveUpdateVACPumpdown is at 500Torr

 The pumpdown has started at 8:38am

Manasa was here to confirm good alignment of the IFO

Inner jam nuts of AC bellow were torqued to 45 ft/lbs and   door nuts were check on opened chambers.

Annulos were roughed down to 500 mTorr

Oplev servos turned off, PSL and green shutters closed before pumpdown started

 

  8969   Tue Aug 6 07:52:58 2013 manasaUpdateGeneralReady to pump down

I did an alignment check of the IFO before we start pumping down.

Arms were locked. PRM and SRM were aligned. Green was aligned to the arms for reference during the pump down.

Steve! It's a GO!

MC spot positions:

MCspot_0806.png

OSEM:

OSEM_0806.png

Oplevs were all centered yesterday and haven't drifted much. So I left them as is.

OL_0806.png

QPDs (IPPOS aligned from yesterday.

QPD_0806.png

  8968   Mon Aug 5 19:10:01 2013 KojiUpdateCDSChoosing LO Amplitudes and Frequencies

- I suppose the green transmission paths were thoroughly inspected and aligned in prior to the measurement

- Of course it is a BAD idea to use 60Hz as the LO frequency.

- Power spectra should be plotted as "RIN (relative intensity noise)" as the DC of 1 and 100 gives you 100 times different power spectra for the same beam.
  Don't forget to subtract the offset from your DC values.

  8967   Mon Aug 5 18:48:44 2013 gautamConfigurationendtable upgradeFull range calibration of PZT mounted mirrors for Y-endtable

 I had prepared two more PZT mounted mirrors for the Y-end some time back. These are:

  • A 2-inch CVI mirror (45 degree, HR and AR for 532nm, was originally one of the steering mirrors at the X-endtable, and was removed while switching those out for the PZT mounted mirrrors).
  • A 1-inch Laseroptik mirror (45 degree, HR and AR for 532nm).

I used the same QPD set-up and the methodology described here to do a full-range calibration of these PZTs. Plots attached. The calibration constants have been determined to be:

CVI-pitch: 0.316 mrad/V

CVI-yaw:  0.4018 mrad/V

Laseroptik pitch: 0.2447 mrad/V

Laseroptik yaw:  0.2822 mrad/V

Remarks:

  • These PZTs, like their X-end counterparts, showed evidence of drift and hysteresis. We just have to deal with this.
  • One of the PZTs (the one on which the CVI mirror is mounted) is a used one. While testing it, I thought that its behaviour was a little anomalous, but the plots do not seem to suggest that anything is amiss.

Plots:

                                                        CVI YAW                                                                                                                         CVI PITCH

2-inch-CVI-Yawcalib.pdf      2-inch-CVI-Pitchcalib.pdf

                                                        Laseroptik YAW                                                                                                             Laseroptik PITCH

1-inch-Laseroptik-Yawcalib.pdf   1-inch-Laseroptik-Pitchcalib.pdf

 

  8966   Mon Aug 5 18:18:32 2013 gautamUpdateCDSChoosing LO Amplitudes and Frequencies

In order to decide what frequencies to dither the 4 degrees of freedom (M1-pitch&yaw, M2-pitch&yaw) at, I took the power spectrum of the X and Y-arm green transmission (C1:ALS-TRX_OUT, C1:ALS-TRY_OUT). Plots showing the power spectra are attached. Looking at the power spectra, I would think that for the X-arm, it would be okay to dither at 40, 50, 60 and 70 Hz. In order to check if the piezos could respond to these frequencies, I used my QPD setup and shook the PZTs with a 100Hz, 1Vpp sinusoid, and saw that the spot moved smoothly on the QPD.


 As for choosing the modulation amplitude, I did a simplistic approximation assuming that the misalignment only rotates the beam axis relative to the cavity axis, and determined what angle coupled 10% of the power into the next eigenmode. Assuming that this is small enough such that if we are already locked to TEM00, the dither won't kick it up to some higher-order mode, the LO amplitude should be in the range of 30-60 digital counts (determined using the PZT calibration constants determined here. This corresponds to a sine-wave of ~50mV amplitude reaching the PZTs (after HV amplification). I am not sure if this is too small, but according to the PZT datasheet, these platforms are supposed to have a resolution of 0.02 urad, which would correspond to the input signal changing by ~0.1 mV, so this signal should be capable of dithering the tip-tilt. 


 I have already added band-pass filters centered at these frequencies to the model (with a passband of 5Hz, 2Hz on either side), and low-pass filters to pull out the DC component of the output of the lock-in amplifiers. It remains to tune the gains of the filter stages. These parameters (frequency, amplitude of the LOs) may also have to be changed after tests). Hopefully the PZTs can be plugged in tomorrow, and I can try and make a measurement of the output matrix. 

Koji also suggested that it may be good to have a path in the model that feeds back to the PZTs by dithering the cavity mirrors as opposed to the PZT mounted mirrors. I will work on incorporating this into the SIMULINK model (c1asx.mdl) and also into the master medm screen.


 

Notes:

  1. The spot size of the X-arm green transmission on the PD was larger than the active surface. I moved the GTRX PD a little back and put in a lens (KPX085, 62.9mm FL, AR.14) in front of the PD, such that the spot is now occupying about 1/4th of the active surface area. The lens was mounted in a Thorlabs LMR1mount, and has been labelled.
  2. I made a slight change to the SIMULINK model, so as to calibrate the PZT sliders to (approximately) volts (I added a multiplier block that multiplies the slider value by constant value 3267.8). The idea is that we can approximately relate the slider value to tilt, knowing the calibration constant in mrad/V for the PZTs.

 

Power Spectra of Arm Green Transmission:

GTR_Power_Spectrum.pdf

  8965   Mon Aug 5 18:02:34 2013 manasaUpdateGeneralClose up list checked
  •  
  • Center beam on all AS optics
  • GET CAMERA IMAGES OF EVERYTHING

    • We must get images right before closing, right after closing, etc.
  • Make sure REFL is clear
    • dither PRM, see motion on AP tables
  • Make sure AS is clear
    • dither BS/ITM, see motion on AP tables
  • Using IPANG/POS pick-off mirrors, center beams on:
    • IPPOS
    • IPANG (aligned low in pitch)
  • Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
  • Check all OpLevs centered, in and out of vacuum

    [Jenne, Manasa]

    IPANG needed to be re-aligned today. Heavy doors are in place and bolts tight (torque 25 & 45).

    Steve! We are ready for pump down!

    I will check the IFO alignment once again early tomorrow morning before Steve starts pumping down.

     

     

     

     

     

  8964   Mon Aug 5 11:53:45 2013 EvanUpdateISSCTN Servo - Explicit Requirement and Proposed Servo

I goofed on the transfer function requirement by not giving you the plant transfer function, which looks to be about 0.014 V/V, independent of frequency (PSL:1278). This needs to be compensated for in the electronic transfer function.

  8963   Mon Aug 5 10:50:48 2013 SteveUpdateVACRGA background

 RGA background at day 12 of this vent . The maglev is pumping on the rga through VM2

 

  8962   Fri Aug 2 22:51:10 2013 JenneUpdateGeneralAll vent tasks complete, just need oplev check

[Manasa, Koji, Jenne]

We went into the BS and IOO chambers, and aligned the green beams such that they came out of the vacuum chamber.  The idea here was to get the beams at the same height, but slightly offset in yaw.  This required moving the Periscope on BS table, PBS in front of that periscope, the Periscope on the IOO table, and 2 steering mirrors on the IOO table after the 2nd periscope.  The tables were not releveled, although we have aligned the full interferometer to this situation, so we do not want to touch the tables.  The MC spot positions are still consistent with those measured earlier this afternoon, before this work, so I'm not concerned.

We confirmed that both green beams are hitting a good place (centered in pitch, and just left and right of center in yaw) on the mirror in the OMC chamber, and are getting to the center of the first mirror on the PSL table.  We then coarsely aligned the beams on the PSL table.

We then relocked and aligned the arms for IR, and checked that the AS beam is centered on the mirrors in the BS chamber, and that the beam is coming out, and to the AS table.  I touched the last mirror before PZT3 a small amount in yaw, and then PZT3 in pitch and yaw, until we saw the beam recentered on the first mirror on the AS table.  At that point, we were also back to the center of the AS camera (which is good, since Koji had aligned all of that the other day).  So, the AS beam is good.

We checked IPPOS, and have centered the beam on all the mirrors, and aligned the beam onto the QPD. 

We checked IPANG, by looking through the viewports at the mirrors in the ETMY chamber.  We are now centered in yaw, but clipping a bit low.  This is what we want, since we always end up drifting high during the pump-down. 

We see a nice, unclipped REFL beam on the camera.

We see a beam a little high on the POP camera, but Koji looked on the table with a card, and saw the beam....we just need to do minor alignment on the out of vac mirrors.

We checked again that the green TEM00 beams from both arms come to the PSL table. 

We are getting POX and POY out, since we are using them to lock and align the arms.

Manasa and Koji recovered one clean allen key from the bottom of the chambers, but one remains, as a sacrifice to the vacuum gods. 

I believe that, with the exception of checking the oplevs and taking photos of PR3, and the green steering optics, we have finished all of our vent tasks.  We should do a quickie alignment on Monday, check the oplevs, take some photos, and put on the heavy doors.  Pumping can start either Monday afternoon or Tuesday morning. 

 

  8961   Fri Aug 2 21:59:36 2013 CharlesUpdateISSFinalized ISS Schematic (hopefully)

Attached is the finalized schematic. The general circuit topology should remain the same from this point forward, although individual component values are subject to change. I will also be adding some more annotations to ensure everything on the board is clear.

In general, I have finally included all of the correct components (i.e. front panel switches are now actually switches and front panel LEDs are now included). I also added an external 'Boost' switch, which can be used to enable or disable the boosts. The motivation for including this switch is that one might want to test functionality of the ISS without using the 'fancy' RMS detection and triggering circuitry. Additionally, one can disable the boosts when all the circuitry is stuffed in order to troubleshoot, so it essentially grants the board some flexibility in its operation.

I am now working on the PCB layout and I should hopefully have that done next week. 

  8960   Fri Aug 2 17:50:10 2013 JenneUpdateGeneralPR3 wedge angle adjusted

[Jenne, Manasa, Koji]

Earlier today, we locked and aligned both the X and Y arms. 

I then went into the BS chamber, put on the BS' aperture, and put an aperture along the AS path.  (We had Michelson fringes, so I centered the aperture around the fringes.  I used one of the brass ruler things that we use to center the beam on ITMs and ETMs, on a riser.  I put this aperture at the edge of the BS table, after the AS beam is launched toward the OMC chamber.  The idea was to replace PR3 such that I could get the beam back through the BS aperture, and the brass ruler aperture, in hopes that we would see arm flashes, and not have to open the ITMY and ETMY heavy doors.)

I set references on the table so that I could put PR3 back in its original position, then removed PR3 from the chamber.

Steve set up a HeNe for me, that we pointed through the optic.  The ghost beam was very high, indicating (as expected) that the wedge was not perfectly horizontal.

I took the suspension off of the cage and laid it down, as I have in the past. 

I removed the optic from the suspension, to try to figure out which was the fat vs. skinny side.  I noticed that there are very faint marks on the actual fat and skinny sides of the optic.  (Mpral - for the LaserOptik mirrors, look for the faint lines that are the full width of the barrel, not the placement of the arrow which marks the HR side).  I put the optic back in (HR side toward the back, fat side on the left (as you look at the face of the optic), which is consistent with the picture in the Optical Layout page of the Wiki, near the bottom.) the optic holder ring.

I put the suspension back on the cage, and saw that the HeNe's ghost beam was now nearly horizontal relative to the straight-through beam.  Excellent.  Also, the pitch balancing didn't seem to change noticably, which I determined was within "poking" distance of where we need it to be.

I put PR3 back onto the BS table, and adjusted it around until I got the beam through both the BS aperture, and the one on the AS path.  As usual, this took quite a while, but as soon as I got through both of those apertures (really at the same place, not close to being through them, but as close as I could tell by eye - this is what took forever), Koji and Manasa saw flashes in the Yarm!  Yay! 

Since I had to move PR3 in angle a tiny bit, I reset the references, then dogged down PR3.  We still had flashes, this time in both arms, so we closed up the light doors.

We have now locked and aligned both arms in IR after the adjustment of PR3, and see both arms' green at 01 or 02. We are about to start checking the green positioning on the periscopes.  We will also need to check the AS path, as well as IPPOS and IPANG before we close up.  We see REFL on the camera.

Separately - Manasa remembered that 2 clean things were dropped yesterday - a screw, and an allen key.  Since they're both Clean, we're not too worried, although she thinks a long-armed person may be able to reach the allen key.

  8959   Thu Aug 1 22:58:45 2013 CharlesUpdateISSCTN Servo - Explicit Requirement and Proposed Servo

 In PSL elog 1270, Evan elucidated the explicit requirements for the CTN ISS board. Essentially, the transfer function of the ISS should be something like:

     TF_mag = (Unstabilized RIN) / (Calculated RIN Requirement)

I took Evan's data and did exactly this. I then designed a servo (using the general design I proposed here) to meet this requirement with a safety factor of ~10. By safety factor, I mean that if the ISS operates exactly according to theory, it should suppress the noise by a factor of 10 more than what is necessary/set out by the requirement. Below is a plot of the loop gain obtained directly from the requirement (the above expression for TF_mag) and the transfer function of the servo I am proposing.

CTN_Servo_TF_-_Proposed_v_Req.png

I don't have the actual schematics attached as I was working with a LISO file and have yet to update the corresponding Altium schematic. The LISO file is attached and I will add the schematics later, although one can reference the second link to find a simple drawing.

  8958   Thu Aug 1 22:49:31 2013 manasaUpdateGeneralGreen status after PR3 flipping investigated

[EricQ, Koji, Manasa]

We opened the BS chamber to check the status of the green beams. The X green has 3 steering mirrors before they hit periscope1 and the Y green transmits through all the optics giving no way to steer it.

We agreed to start fixing the Y green. The wedge angle of PR3 is steering the transmitted beam away in both pitch and yaw. Since we are restricted only to yaw movement (done by moving the periscope), we want he wedge angle to be oriented in the yaw as well.

Right now, the wedge is oriented at about 20-30 deg off (The mark on the side of the mirror does not indicate the wedge). So we see a pitch as well as yaw misalignment in the transmitted beam. The pitch misalignment is making the beam fall off the mirrors in periscope2. 

We have decided to get the wedge angle set right for PR3 and redo the alignment for IR. Once we are aligned for the IR, we will modify the green layout.

  8957   Thu Aug 1 21:28:09 2013 gautamUpdateCDSSlow channels set-up in ALS

The following slow channels have been added and are now being recorded by FB.

 

C1:ALS-X_OVEN_TEMP

C1:ALS-Y_OVEN_TEMP

C1:ALS-BEATX_FREQ

C1:ALS-BEATY_FREQ


Details:

In order to integrate the data collected by the Raspberry-Pi from the Y-end doubling oven temperature controller and also the data from the frequency counter which will be hooked up to monitor the beat frequency, Koji helped me set up some slow EPICS record channels (in ALS as we felt this was most appropriate). The procedure for setting up slow channels was as follows (virtually identical to what is detailed in this elog:

  1. Add the channel names to the file C0EDCU.ini (path = /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini).
  2. Make a database (.db) file so that these channels are actually recorded (path = /cvs/cds/caltech/target/c1aux/als.db).
  3. Restart framebuilder. 
  4. Verify that the channels indeed exist and can be read and written to using ezcaread and ezcawrite.

I will now integrate these channels into my scripts, and make some simple MEDM screens.

 

  8956   Thu Aug 1 20:58:56 2013 gautamUpdateCDSNew model for endtable PZTs-MEDM Screens setup

 

I have made some minor changes to the model, made all the MEDM screens, and linked monitors on these to the appropriate channels. I have borrowed heavily from the C1ASS MEDM screens (particularly for the small filter modules-it was convenient to just copy and paste an existing module, and edit the channel names using EMACS/GEDIT), and have edited these to suit the needs of this servo. Some features:

  • The feedback signal (only the output of the servo to the PZTs, plus any contribution from the on-screen sliders, and not including the LO output) is monitored with both a slow (using CDS_EPICS_OUTPUT block from the CDS_PARTS library) and fast channel (using Test Point from the same library). The idea is that it would be useful to know the output to the PZTs such that if coarse adjustment ever needs to be done at the endtable, the PZTs can be restored to the middle of its operating range by means of the sliders.
  • Sliders are incorporated into the master screen for adjusting the output to the PZTs. There are text-input fields below the sliders as well, which control the same channel.
  • I have removed the 4 remaining excitation points to the DAC set up in C1SCX, and have relocated them to channels 12-15 of the DAC in C1ASX.

I think I am now ready to take some measurements and try and optimize this servo. There is no green transmission at the PSL table at the moment, so not much can be done, though the first step would be to take the power spectrum of the error signal, which would help me decide the appropriate frequencies for the LOs. I would then have to add the appropriate filters to the model. The last, and most difficult step, would be the measurement of the output matrix, though Koji has given me some ideas about how this measurement can be done. I also have a template script ready, though I will only finalise this after optimising the servo and running it a couple of times manually.

 

Attached are screenshots of the MEDM screens.

 

MAIN_SCREEN.pdf      MATRICES.pdf   

LOCKINS.pdf      CONTROL_FILTERS.pdf

 

  8955   Thu Aug 1 18:55:20 2013 Alex ColeUpdateElectronicsPreliminary Photodetector Frequency Reponse Measurements

Quote:

[Eric, Alex]

We used our setup from yesterday (elog #8940) to measure transimpedance measurements for AS55, REFL11, REFL33, and REFL55, using our Newport 1611 FC-AC as reference. We connected the fibers to their respective telescopes such that the beams focused on their photodetectors, using a multimeter to maximize photodetector DC output. Plots are attached. At first glance, the poles seem to be where they're supposed to be.

Note that the procedure used today is similar to what the eventual automated procedure will be. The main differences are (1) The RF Switch will be used rather than manual switching (2) NWAG4395A will be used to collect data rather than netgpibdata (3) Data will be fit using vectfit4.m and compared to some canonical set.

 [Alex, Eric]

Today I spent some time mounting the launcher and performing the same data collection for POX11. I think I still need to focus the launcher so the photodetector gets a good signal, but the data from today wasn't too bad.  Additionally, I worked on matlab scripts to improve PDFR data analysis.

This time I collected data from the network analyzer using NWAG4395A in the netgpibdata directory. The advantage of this is that the computer tells the network analyzer to perform the sweep as well as retrieving the data.

For analysis, I improved my implementation of vectfit4.m so that it focuses in on the particular photodetector's predicted peaks and thus ignores much of the noise, giving a better fit. The raw data is the red circles in the 2nd attachment, while the fit is the blue line. I also had the program return the frequency value of the peak. For POX11, this was 1.106e+07 Hz.

I also finagled copies of existing programs to enable one to plot multiple transfer functions on the same axes. This function is /users/alex.cole/plottwo.m. I will eventually use this to compare new data to some canonical data so that we may monitor photodetector performance over time.

The eventual plan is to generate two plots per photodetector, one of which will compare new data to the canonical set, the other of which will show the fit of the data. Both will have subplots that zoom in around regions of interest (known peaks and notches), and the plot which displays the canonical set will also have Q's of peaks and their locations.

  8954   Thu Aug 1 16:33:39 2013 SujanUpdatePEMGuralp2 seismometer installed at North side of POX table

1)Power to the seismometers were turned down,

2)Guralp2 was moved to North side of POX table

3)Guralp2 was aligned in N-s Direction and leveled before connecting

4)Power to seismometers was turned on once Guralp2 was connected

IMG_1213.JPG

  8953   Thu Aug 1 16:03:52 2013 SteveUpdateGeneralreopen BS & IOO

Quote:

[Koji, Manasa]

We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.

We should fix this tomorrow. The BS heavy door should come off.

Steve! Do not start pump down tomorrow !

 We closed the chambers last night with heavy doors and reopened it today.

  8952   Thu Aug 1 15:28:44 2013 gautamUpdateCDSNew model for endtable PZTs-problem solved

Quote:

 

I don't know what's going on here (why the channels are white), and I don't yet have a suggestion of where to look to fix it but...

Is there a reason that you're making a new model for this?  You could just use and existing model at c1iscex, like the c1scx, and put your stuff in a top-names block.  Then you wouldn't have to worry about all of the issues with adding and integrating a new model.

Koji just fixed this.

It seems that the new model's channels were not automatically added to the master file in the framebuilder (/opt/rtcds/caltech/c1/target/master). Adding the following two lines to the master file fixed the problem;

/opt/rtcds/caltech/c1/chans/daq/C1ASX.ini

/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1asx.par

The box is now green. It looks like C1ASX.ini is created automatically in /opt/rtcds/caltech/c1/chans but the master file needs to be manually edited. The channels are now showing up on dataviewer etc. I have updated the information on the wiki page.


 

 

 

 

 

  8951   Thu Aug 1 15:06:59 2013 jamieUpdateCDSNew model for endtable PZTs

Quote:

I have made a new model for the endtable PZT servo, and have put it in c1iscex. Model name is c1asx. Yesterday, Koji helped me start the model up. The model seems to be running fine now (there were some problems initially, I will post a more detailed elog about this in a bit) but some channels, which are computer generated, don't seem to exist (they show up as white blocks on the MEDM GDS_TP screen). I am attaching a screenshot of the said screen and the names of the channels. More detailed elog about what was done in making the model to follow.

 

C1ASX_GDS_TP.png

 

Channel Names:

C1:DAQ-DC0_C1ASX_STATUS (this is the channel name for the two leftmost white blocks)

C1:DAQ_DC0_C1ASX_CRC_CPS

C1:DAQ-DC0_C1ASX_CRC_SUM

I don't know what's going on here (why the channels are white), and I don't yet have a suggestion of where to look to fix it but...

Is there a reason that you're making a new model for this?  You could just use and existing model at c1iscex, like the c1scx, and put your stuff in a top-names block.  Then you wouldn't have to worry about all of the issues with adding and integrating a new model.

  8950   Thu Aug 1 13:09:17 2013 gautamUpdateCDSNew model for endtable PZTs-procedure

 

 These are roughly the steps I followed in setting up the new model for the endtable PZT servo - C1ASX.


Simulink model:

I made a SIMULINK model of the servo, using MATLAB R2013a. The path to the model is /opt/rtcds/caltech/c1/userapps/release/isc/c1/models/c1asx.mdl. I am listing the parameters set on the CDS_PARAMETERS block:

  • host = c1iscex
  • site = c1
  • rate = 16k
  • dcuid = 44 (which I chose after making sure that this dcuid was not used on this list which was last updated end Feb 2013)
  • specific_cpu = 5 (again chosen after checking the available CPUs in the above list).
  • adc_Slave = 1
  • shmem_daq = 1
  • no_rfm_dma = 1
  • biquad = 1

 

Making, Compiling and Installing the Model:

After saving the model, I ssh-ed into c1iscex and ran the following commands:

rtcds make c1asx - this gave me a whole bunch of errors initially, which I tracked down to a naming problem in some of the from and goto flags: there should not be any spaces.

rtcds install c1asx 

rtcds start c1asx - this gave me an error which said something like 'can't start/stop model.' Koji pointed out that given that a new model is being started, there is an additional step involved, which is to add the model name to the rtsystab file (this is located at /diskless/root/etc/rtsystab on framebuilder, and is mirrored in the various computers. It would be advisable to make sure that the changes are mirrored in the corresponding file on the computer in which the new model is being installed). 

After adding the model name to the rtsystab file, I tried running rtcds start c1asx again. This time, no errors were output, but the model was not up and running as verified by looking at the C1:ASX_GDS_TP medm screen.


 Debugging 

Koji suggested making a simple model (1 CDS parameters block, 1 ADC block and 2 filter modules, appropriately terminated) and see if that starts up, which it did. I then tried adding my servo minus the DAC block and recompiled and restarted the model. This too worked fine. I figured that the next logical step would be to add the DAC block to the model, and restart the model. But when I tried this, c1iscex crashed .

Jenne helped in restoring things to a working state (we reverted the c1asx model to just 2 filter modules, and went to the X-end and restarted the computer. This did not work the first time so I went back in and restarted it again, at which point we were able to ssh into c1iscex again and restart the four models running on it).

Since Manasa and Koji were working on getting things set up for the pumpdown,I did not try anything again till later in the evening, when Koji helped in debugging the problem further. In the meantime, at Jenne' suggestion, I made the model once again in MATLAB R2010b. In the evening, when I tried restarting the model, Koji suggested that the DAC channels in c1asx may be used by other models, at which point I realised I had set up excitation points on channels 8 through 15 of the DAC in c1scx (detailed here) in order to test the hardware at 1X9. I removed the excitation points from channels 8-11 of the DAC block in c1scx (these are the ones used in c1asx), and recompiled and restarted c1asx (using the above sequence of commands). I then tried recompiling and starting c1asx once more, and this time, it worked . At least, the GDS_TP screen suggests that the model is running alright, except for the fact that some computer generated channels seem to be missing. This problem is unresolved for now, and probably has something to do with the fact that C1ASX channels do not appear in Dataviewer.

I do not think this has to do with restarting framebuilder (I did the usual telnel fb 8088 followed by shutdown). In any case, I have added the new model to the CDS_FE_STATUS screen, and will continue to debug the same. I have also got a template medm screen (work in progress) which I will elog about soon as I get it done.

 

Note to self: There are 4 more excitation channels still hooked up to the DAC (channels 12-15) in the c1scx model. I plan to remove these and put them in c1asx.

 

  8949   Thu Aug 1 12:12:35 2013 gautamUpdateCDSNew model for endtable PZTs

I have made a new model for the endtable PZT servo, and have put it in c1iscex. Model name is c1asx. Yesterday, Koji helped me start the model up. The model seems to be running fine now (there were some problems initially, I will post a more detailed elog about this in a bit) but some channels, which are computer generated, don't seem to exist (they show up as white blocks on the MEDM GDS_TP screen). I am attaching a screenshot of the said screen and the names of the channels. More detailed elog about what was done in making the model to follow.

 

C1ASX_GDS_TP.png

 

Channel Names:

C1:DAQ-DC0_C1ASX_STATUS (this is the channel name for the two leftmost white blocks)

C1:DAQ_DC0_C1ASX_CRC_CPS

C1:DAQ-DC0_C1ASX_CRC_SUM

  8948   Wed Jul 31 21:12:05 2013 manasaUpdateGeneralPump down called off

[Koji, Manasa]

We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.

We should fix this tomorrow. The BS heavy door should come off.

Steve! Do not start pump down tomorrow !

  8947   Wed Jul 31 17:02:17 2013 Alex ColeUpdateElectronicsPreliminary Photodetector Frequency Reponse Measurements

[Eric, Alex]

We used our setup from yesterday (elog #8940) to measure transimpedance measurements for AS55, REFL11, REFL33, and REFL55, using our Newport 1611 FC-AC as reference. We connected the fibers to their respective telescopes such that the beams focused on their photodetectors, using a multimeter to maximize photodetector DC output. Plots are attached. At first glance, the poles seem to be where they're supposed to be.

Note that the procedure used today is similar to what the eventual automated procedure will be. The main differences are (1) The RF Switch will be used rather than manual switching (2) NWAG4395A will be used to collect data rather than netgpibdata (3) Data will be fit using vectfit4.m and compared to some canonical set.

  8946   Wed Jul 31 10:27:00 2013 SteveUpdateVACinstrument grade air is discontinoued

We have been using Air Liquide Instrument Grade Air for venting the 40m vacuum envelope. It is no longer available.

The replacement is Alphagas 1, total hydrocarbon <0.1% ppm

LINK TO Alphagaz

 

A- A+ Imprimer Partager

ALPHAGAZ™ Specifications

ALPHAGAZ™ range of Pure Gases for Analysis: two purity levels tailored to cover all your analytical needs.

Because each analytical application is sensitive to critical impurities that can affect your results, tests in our research center and analyzer manufacturers' studies confirm that the main impurities are moisture, oxygen and hydrocarbons.

Based on these critical impurities, Air Liquide has developed ALPHAGAZ™ brand with a Maximum Impurity Level (H2O, O2, CnHm).

  • The Maximum Impurity Level is the same across the whole product range. This constitutes a guarantee of quality with regard to major critical impurities and contamination tracers. ALPHAGAZ™ 2 offers an even greater guarantee in terms of purity (CO, CO2, H2).
  • The Minimum Total Purity is a prerequisite for product quality, but we do not stop there: Air Liquide selects and controls its gas sources by defining specifications regarding impurities that are critical for your analysis or application.

 

ALPHAGAZ™ 1 ALPHAGAZ™ 1 ALPHAGAZ™ 2 ALPHAGAZ™ 2 ALPHAGAZ™ AUTO IV ALPHAGAZ™ CO2 SFC
He, N2, H2,
Ar, O2
Air He, N2, H2,
Ar, O2
Air Air*, N2, O2 CO2
Specifications (ppm vol):
H2O < 3 ppm
O2 < 2 ppm
CnHm < 0.5 ppm
H2O< 3 ppm
CnHm< 0.1 ppm
CO2 < 1 ppm
CO < 1 ppm
H2O < 0.5 ppm
O2 < 0.1 ppm
CO2 < 0.1 ppm
CO < 0.1 ppm
CnHm < 0.1 ppm
H2 < 0.1 ppm
H2O < 0,5 ppm
CnHm < 50 ppb
CO2 < 0,1 ppm
CO < 0,1 ppm
NOx < 10 ppb *
SO2 < 10 ppb*
CnHm < 0.2 ppm
CO2 < 0.5 ppm
CO < 0.5 ppm
NO/NOx < 0.1 ppm
H2O < 5 ppm
O2 < 2 ppm
CnHm < 2 ppm
NVOC* < 2 ppm
CO < 5 ppm
NO + NO2 < 2 ppm
S Total < 1 ppm
H2 < 0.5 ppm
  20 ± 1 % O2 (Europe)
21 ± 1% O2 (North America)
  20,9 ± 1 % O2
*: Detection limit
*20 ± 1 % O2 *NVOC: Non Volatile Organic Compounds
Minimum Total Purity

 

 

ELOG V3.1.3-